OpenStudy (anonymous):

solve:-(3x-2)=-4(x+1)-4

OpenStudy (debbieg):

Distribute through the parens fist... what do you get on each side?

OpenStudy (anonymous):

i got 2

OpenStudy (anonymous):

Alrigh pana here we go:\[-3x+2=-4x-4-4\]\[-3x+4x=-2-4-4\]\[x=-10\]

OpenStudy (anonymous):

i cant see that

OpenStudy (anonymous):

-3x+2=-4x-4-4 -3x+4x=-2-4-4 x=-10 Better domi?

OpenStudy (anonymous):

yes thanks :)